OpenStudy (anonymous):

3. Segment AB is the diameter of circle M. The coordinates of A are (−4,3). The coordinates of M are (1,5). What are the coordinates of B? (1) (6,7) (3) (−3,8) (2) (5,8) (4) (−5,2)

OpenStudy (mertsj):

To get from A (-4,3) to M (1,5) you must go 5 to the right and up 2. Do that one more time and you will be at the other end of the diameter or go right 10 and up 4 from the original point and arrive at (6,7).

OpenStudy (anonymous):

how do i show work????

OpenStudy (mertsj):

you could say m, the slope of AM = (3-5)/(-4-1)= -2/-5 = 2/5 Therefore the slope of MB must be the same so from Point M, if I rise 2 and run 5 I will be at (6.7)